# EXAONE 4.0: Unified Large Language Models Integrating Non-reasoning and Reasoning Modes

### v1

@@ 1 Introduction
The previous version, EXAONE 3.5 [31], focused on real-world usability by strengthening comprehensive instruction-following abilities, while EXAONE Deep [32] emphasized reasoning performance, particularly in mathematical and coding domains.
@@ 2.1 Model Configurations
Contrary to past findings that models using global attention across all layers performed better, recent studies [14, 15, 36] suggested that utilizing a larger window size (e.g., from 512 to 1,024 or 4,096) and applying global attention to only a minority of layers can still achieve excellent long-context performance. Additionally, it reported that incorporating small amounts of global attention periodically in combinations with heterogeneous structures like Mamba [43, 34, 39] helped maintain the ability to understand global context.
In designing the EXAONE 4.0 model, a sliding window size of 4K is selected to minimize any adverse effects on short-context performance. Furthermore, the model does not employ Rotary Position embedding [63] for global attention, ensuring that the model does not develop biases towards length and can maintain a global view. For the design of the local attention mechanism, we do not employ the chunked attention strategy. Instead, we adopt sliding window attention, a well-established form of sparse attention that offers strong theoretical stability. Unlike chunked attention, sliding window attention benefits from wide support in open-source frameworks, ensuring robust implementation and ease of integration. To prevent performance degradation in short-context areas during long context fine-tuning, the EXAONE 4.0 model employs a careful data selection methodology and a progressive training recipe, effectively balancing efficiency and performance.
Another significant change in the EXAONE 4.0 model is the repositioning of layer normalization (LayerNorm), as shown in Figure 2. According to recent studies [53], some layers that do not significantly impact model performance are found mainly in deep layers. This issue is attributed to the Pre-LN transformer architecture [61], which enhances stability but leads to exponentially increasing variance in outputs as model depth increases. A simple operation to control variance by providing more scaling to outputs as layers deepen was proposed, but we find that the QK-Reorder-LN method [56, 42], which applies LayerNorm after input queries and keys and performs LayerNorm after attention output, yields better performance on downstream tasks despite consuming more computation. The normalization type RMSNorm, which was applied since EXAONE 3.0, is retained in EXAONE 4.0.
@@
Non-linearity SwiGLU [50] SwiGLU
@@
Head type GQA [4] GQA
@@
RoPE theta [52] 1,000,000 1,000,000
Tokenizer BBPE [58] BBPE
@@ 2.2 Pre-training
The amount of data and computational resources used for pretraining in the EXAONE 4.0 models are summarized in Table 2. For the EXAONE 3.5 32B model, 6.5 trillion tokens are used for pretraining. In comparison, the EXAONE 4.0 32B model doubles this amount, utilizing 14 trillion tokens for pretraining. This increase in data is specifically aimed at enhancing the model’s world knowledge. As will be discussed later, this approach yields noticeable improvements in benchmarks that rely on knowledge, such as MMLU-Redux [13], where the use of more extensive training data has a demonstrable impact on performance.
@@
Furthermore, as recent studies showed that reasoning performance was significantly influenced by the cognitive behavior [12] acquired from documents seen during pretraining, we perform rigorous data curation during pretraining to enhance post-training performance.
@@ 2.3 Context Length Extension
The long-context fine-tuning process is meticulously executed, with the Needle In A Haystack (NIAH) test [16] at each stage to ensure thorough validation of the model’s performance. This iterative refinement continues until comprehensive optimization is achieved and the “green light” signal is consistently observed across all segments, signifying the successful extension of the context length to 128K tokens without compromising the model’s overall performance.
@@ 2.4.2 Reasoning Reinforcement Learning
To enhance the model’s reasoning capabilities, we conduct online reinforcement learning (RL) following supervised fine-tuning (SFT). Previous studies demonstrated that combining the GRPO (Group Relative Policy Optimization) algorithm [49] with verifiable rewards [28] can effectively improve model performance. To address the limitations of existing GRPO, we propose a new algorithm, named AGAPO (Asymmetric Sampling and Global Advantage Policy Optimization).
@@
  • Remove Clipped Objective. Previous research has questioned the necessity of PPO(Proximal Policy Optimization) [48]’s clip loss [3] and shown that this clipped objective can degrade performance [40] by preventing crucial, low-probability tokens from contributing to gradient updates. These tokens are often associated with reflective behaviors that serve as forks in the reasoning path. AGAPO removes the clipping from PPO and instead uses a standard policy gradient loss. This approach is designed to prevent the dropping of these exploratory tokens, allowing for more substantial policy updates while maintaining training stability.
  • Asymmetric Sampling. Previous works filter out samples where all responses were either correct or incorrect [67, 17] because they result in a zero advantage for GRPO. However, as recent work has shown the effectiveness of Negative Sample Reinforcement [70], AGAPO utilizes an asymmetric sampling method that does not discard samples where all responses are incorrect, thereby including a higher proportion of negative feedback. For these all-incorrect samples, a small negative reward is assigned through the advantage calculation, allowing them to be used to guide the model away from erroneous reasoning paths.
  • Group&Global Advantages. GRPO advantage method does not account for the distribution of the entire batch, which makes it difficult to assign appropriate negative rewards to groups of all-incorrect samples. To improve this, AGAPO calculates the advantage in two stages: group and global. First, at the group level, the advantage is computed using the Leave-One-Out(LOO) method [3] within a response group based on verifiable reward accuracy. Next, normalization is performed across the entire batch (global) to calculate a final advantage that considers the full batch distribution.
  • Sequence Level Cumulative KL. To enhance performance while preserving capabilities learned during the SFT stage, a KL penalty is applied. We adopt the sequence-level cumulative KL [54], as proposed in prior research, to ensure the model receives an appropriate gradient during training.
@@ 2.4.3 Preference Learning with Hybrid Reward
Preference learning is conducted by directly learning human preferences from chosen and rejected data pairs, akin to the Direct Policy Optimization (DPO) framework [46]. We employ SimPER [60], among various reference-free preference optimization methods for this learning process. The dataset for preference learning is constructed based on on-policy responses [38, 28] generated by the model after completing the RL phase. For each query, we generate 4 to 16 responses per task, and select chosen and rejected responses based on a hybrid reward combining verifiable reward, preference reward, language consistency reward, and conciseness reward, tailored per task.
@@ 2.5 Data Compliance
Developing AI models requires a large amount of data, and the acquisition and utilization of this data can lead to various legal issues, such as copyright infringement, intellectual property infringement, and personal information protection violations. To minimize these risks, LG AI Research conducts AI Compliance reviews throughout the entire process of data collection, AI model training, and information provision. For more detailed information, please refer to the EXAONE 3.0 Technical Report [30] and the LG AI Ethics Principles [29].
@@ 3.1 Benchmarks
  • World Knowledge We select benchmarks to evaluate the extent of our model’s world knowledge, including MMLU-Redux [13] and MMLU-Pro [59], a refined and extended version of MMLU [19]. Additionally, we utilize GPQA-Diamond [47] to assess the expert-level knowledge in Biology, Physics, and Chemistry.
@@
  • Math/Coding Challenging benchmarks in Math and Coding categories are adopted to evaluate the test-time computational capability of EXAONE 4.0. For Math, we utilize two math Olympiad competitions: AIME 2025 [37] and HMMT Feb 2025 [6]. For Coding, LiveCodeBench v5 and v6 [24] are chosen.
@@
  • Instruction Following To evaluate how well our models understand and align with users’ instructions, we select IFEval [69] and Multi-IF [18], the latter being an extension of IFEval to support multi-turn and multilingual scenarios. We use only the English subset of Multi-IF to focus on assessing the multi-turn instruction-following ability on English.
@@
  • Long Context To evaluate the model’s ability to understand and solve tasks requiring long-context comprehension, we adopt three representative benchmarks: HELMET [66], RULER [22], and LongBench [5]. These benchmarks collectively cover both synthetic tasks and real-world scenarios. To maintain a coherent evaluation focused on core long-context capabilities, we exclude the LongCite task from HELMET (see Appendix D for further details).
@@
  • Agentic Tool Use With the advancement of LLM-based agents, numerous benchmarks have emerged to evaluate their tool-use capabilities, among which we focus on the two most widely adopted: BFCL-v3 [44] and Tau-Bench [65]. BFCL-v3 evaluates various aspects of function-calling abilities. Tau-Bench assesses tool calling performance through simulated conversations between a user LLM. We utilize gpt-4.1-2025-04-14 model as the user role.
@@
  • Multilinguality Beyond English, we evaluate our models on two additional languages: Korean and Spanish. For Korean, we use KMMLU-Pro [^0] https://huggingface.co/datasets/LGAI-EXAONE/KMMLU-Pro for measuring practical applicability on professional knowledge and KMMLU-Redux [^1] https://huggingface.co/datasets/LGAI-EXAONE/KMMLU-Redux for assessing real-world expert knowledge instead of KMMLU [51] to ensure benchmark reliability. KMMLU have been reported dataset error and contamination issue between pre-training corpora and task dataset. In addition, we employ Korean School Math (KSM) subset of HRM8K [26] to evaluate a wide range of Korean mathematical knowledge from high-school to Olympiad level. To evaluate the models’ ability to handle long-context Korean inputs, we also include an in-house benchmark, Ko-LongBench (Please refer to Appendix D.4 for details). For Spanish, we adopt the translated version of existing benchmarks. To be specific, we use MMMLU (es)  [^2] https://huggingface.co/datasets/openai/MMMLU and MATH500 [35] (es)  [^3] https://huggingface.co/datasets/bezir/MATH-500-multilingual. Furthermore, we assess translation ability using WMT24++ [10], a widely-used translation benchmark. We consider only EN and ES pair, and utilize LLM-as-a-judge 444gpt-4.1-2025-04-14 is used for the judge model. We follow reference-based direct assessment method used in WMT24++ [10]. The exact prompt used for judge is in Appendix D.5. to score the translation quality.
@@ 3.3 Experimental Setup
We sample $n$ different responses for each problem in benchmarks with limited examples to ensure evaluation stability. Specifically, we sample $n=8$ responses for GPQA-Diamond, $n=32$ for AIME 2025 and HMMT Feb 2025, and $n=4$ for LiveCodeBench v5/6, Tau-Bench and MATH500 (es). The accuracy is averaged over the $n$ samples. In Reasoning mode, we set temperature to 0.6, top-p [21] to 0.95, and apply a presence penalty of 1.5 only for our 32B model. In contrast, for Non-Reasoning mode, greedy decoding is used for a single ($n=1$) generated response, while the same sampling settings as Reasoning mode (except with a presence penalty of 0.0) are used when generating $n>1$ responses. We generate a maximum of 64K tokens for AIME 2025, HMMT Feb 2025, LiveCodeBench v5/6, and KSM benchmarks, while 32K for other benchmarks.
@@
In evaluating long-context performance of Small-Size NON-REASONING models, we extend the context lengths of Qwen3 1.7B and Qwen3 0.6B beyond their 32K token limit by applying YaRN [45], enabling inference up to 64K tokens. For reference, evaluation results of models such as Gemma-3-1B, EXAONE-3.5-2.4B-Instruct, Qwen3 1.7B, and Qwen3 0.6B at context lengths up to 32K tokens are provided in the Appendix D.
@@ Appendix C Baseline models
Frontier DeepSeek R1-0528 [8] 671B (MoE) Reasoning May 2025
DeepSeek V3-0324 [9] 671B (MoE) Non-reasoning Mar. 2025
@@
Qwen 3 235B [62] 235B (MoE) Hybrid Apr. 2025
Mid-size Qwen 3 32B [62] 32.8B Hybrid Apr. 2025
@@
Gemma 3 27B [55] 27.4B Non-reasoning Mar. 2025
@@
Magistral-Small-2506 [41] 23.6B Reasoning Jun. 2025
Phi 4 reasoning plus [1] 14.7B Reasoning Apr. 2025
Phi 4 [2] 14.7B Non-reasoning Dec. 2024
@@
EXAONE Deep 2.4B [32] 2.41B Reasoning Mar. 2025
Qwen 3 1.7B [62] 1.72B Hybrid Apr. 2025
@@
Gemma 3 1B [55] 1.00B Non-reasoning Mar. 2025
Qwen 3 0.6B [62] 596M Hybrid Apr. 2025

### v2

@@ 1 Introduction
The previous version, EXAONE 3.5 [research2024exaone35serieslarge], focused on real-world usability by strengthening comprehensive instruction-following abilities, while EXAONE Deep [research2025exaonedeepreasoningenhanced] emphasized reasoning performance, particularly in mathematical and coding domains.
@@ 2.1 Model Configurations
Contrary to past findings that models using global attention across all layers performed better, recent studies [gemma2, gemma3, llama4] suggested that utilizing a larger window size (e.g., from 512 to 1,024 or 4,096) and applying global attention to only a minority of layers can still achieve excellent long-context performance. Additionally, it reported that incorporating small amounts of global attention periodically in combinations with heterogeneous structures like Mamba [nvidia2025nemotronhfamilyaccurateefficient, lieber2024jambahybridtransformermambalanguage, minimax2025minimax01scalingfoundationmodels] helped maintain the ability to understand global context.
In designing the EXAONE 4.0 model, a sliding window size of 4K is selected to minimize any adverse effects on short-context performance. Furthermore, the model does not employ Rotary Position embedding [yang2025ropenopeagainnew] for global attention, ensuring that the model does not develop biases towards length and can maintain a global view. For the design of the local attention mechanism, we do not employ the chunked attention strategy. Instead, we adopt sliding window attention, a well‑established form of sparse attention that offers strong theoretical stability. Unlike chunked attention, sliding window attention benefits from wide support in open‑source frameworks, ensuring robust implementation and ease of integration. To prevent performance degradation in short-context areas during long context fine-tuning, the EXAONE 4.0 model employs a careful data selection methodology and a progressive training recipe, effectively balancing efficiency and performance.
Another significant change in the EXAONE 4.0 model is the repositioning of layer normalization (LayerNorm), as shown in Figure 2. According to recent studies [sun2025cursedepthlargelanguage], some layers that do not significantly impact model performance are found mainly in deep layers. This issue is attributed to the Pre-LN transformer architecture [xiong2020layernormalizationtransformerarchitecture], which enhances stability but leads to exponentially increasing variance in outputs as model depth increases. A simple operation to control variance by providing more scaling to outputs as layers deepen was proposed, but we find that the QK-Reorder-LN method [olmo20252olmo2furious, muennighoff2025olmoeopenmixtureofexpertslanguage], which applies LayerNorm after input queries and keys and performs LayerNorm after attention output, yields better performance on downstream tasks despite consuming more computation. The normalization type RMSNorm, which was applied since EXAONE 3.0, is retained in EXAONE 4.0.
@@
Non-linearity SwiGLU [shazeer2020gluvariantsimprovetransformer] SwiGLU
@@
Head type GQA [ainslie2023gqatraininggeneralizedmultiquery] GQA
@@
RoPE theta [su2023roformerenhancedtransformerrotary] 1,000,000 1,000,000
Tokenizer BBPE [wang2020bbpe] BBPE
@@ 2.2 Pre-training
The amount of data and computational resources used for pretraining in the EXAONE 4.0 models are summarized in Table 2. For the EXAONE 3.5 32B model, 6.5 trillion tokens are used for pretraining. In comparison, the EXAONE 4.0 32B model doubles this amount, utilizing 14 trillion tokens for pretraining. This increase in data is specifically aimed at enhancing the model’s world knowledge. As will be discussed later, this approach yields noticeable improvements in benchmarks that rely on knowledge, such as MMLU-Redux [gema2025mmlu], where the use of more extensive training data has a demonstrable impact on performance.
@@
Furthermore, as recent studies showed that reasoning performance was significantly influenced by the cognitive behavior [gandhi2025cognitivebehaviorsenableselfimproving] acquired from documents seen during pretraining, we perform rigorous data curation during pretraining to enhance post-training performance.
@@ 2.3 Context Length Extension
The long-context fine-tuning process is meticulously executed, with the Needle In A Haystack (NIAH) test [niah] at each stage to ensure thorough validation of the model’s performance. This iterative refinement continues until comprehensive optimization is achieved and the “green light” signal is consistently observed across all segments, signifying the successful extension of the context length to 128K tokens without compromising the model’s overall performance.
@@ 2.4.2 Reasoning Reinforcement Learning
To enhance the model’s reasoning capabilities, we conduct online reinforcement learning (RL) following supervised fine-tuning (SFT). Previous studies demonstrated that combining the GRPO (Group Relative Policy Optimization) algorithm [shao2024deepseekmathpushinglimitsmathematical] with verifiable rewards [lambert2025tulu3pushingfrontiers] can effectively improve model performance. To address the limitations of existing GRPO, we propose a new algorithm, named AGAPO (Asymmetric Sampling and Global Advantage Policy Optimization).
@@
  • Remove Clipped Objective. Previous research has questioned the necessity of PPO(Proximal Policy Optimization) [schulman2017proximalpolicyoptimizationalgorithms]’s clip loss [ahmadian2024basicsrevisitingreinforcestyle] and shown that this clipped objective can degrade performance [minimax2025minimaxm1scalingtesttimecompute] by preventing crucial, low-probability tokens from contributing to gradient updates. These tokens are often associated with reflective behaviors that serve as forks in the reasoning path. AGAPO removes the clipping from PPO and instead uses a standard policy gradient loss. This approach is designed to prevent the dropping of these exploratory tokens, allowing for more substantial policy updates while maintaining training stability.
  • Asymmetric Sampling. Previous works filter out samples where all responses were either correct or incorrect [yu2025dapoopensourcellmreinforcement, he2025skyworkopenreasoner1] because they result in a zero advantage for GRPO. However, as recent work has shown the effectiveness of Negative Sample Reinforcement [zhu2025surprisingeffectivenessnegativereinforcement], AGAPO utilizes an asymmetric sampling method that does not discard samples where all responses are incorrect, thereby including a higher proportion of negative feedback. For these all-incorrect samples, a small negative reward is assigned through the advantage calculation, allowing them to be used to guide the model away from erroneous reasoning paths.
  • Group&Global Advantages. GRPO advantage method does not account for the distribution of the entire batch, which makes it difficult to assign appropriate negative rewards to groups of all-incorrect samples. To improve this, AGAPO calculates the advantage in two stages: group and global. First, at the group level, the advantage is computed using the Leave-One-Out(LOO) method [ahmadian2024basicsrevisitingreinforcestyle] within a response group based on verifiable reward accuracy. Next, normalization is performed across the entire batch (global) to calculate a final advantage that considers the full batch distribution.
  • Sequence Level Cumulative KL. To enhance performance while preserving capabilities learned during the SFT stage, a KL penalty is applied. We adopt the sequence-level cumulative KL [tang2025pitfallskldivergencegradient], as proposed in prior research, to ensure the model receives an appropriate gradient during training.
@@ 2.4.3 Preference Learning with Hybrid Reward
Preference learning is conducted by directly learning human preferences from chosen and rejected data pairs, akin to the Direct Policy Optimization (DPO) framework [rafailov2024directpreferenceoptimizationlanguage]. We employ SimPER [xiao2025simperminimalistapproachpreference], among various reference-free preference optimization methods for this learning process. The dataset for preference learning is constructed based on on-policy responses [meng2024simposimplepreferenceoptimization, lambert2025tulu3pushingfrontiers] generated by the model after completing the RL phase. For each query, we generate 4 to 16 responses per task, and select chosen and rejected responses based on a hybrid reward combining verifiable reward, preference reward, language consistency reward, and conciseness reward, tailored per task.
@@ 2.5 Data Compliance
Developing AI models requires a large amount of data, and the acquisition and utilization of this data can lead to various legal issues, such as copyright infringement, intellectual property infringement, and personal information protection violations. To minimize these risks, LG AI Research conducts AI Compliance reviews throughout the entire process of data collection, AI model training, and information provision. For more detailed information, please refer to the EXAONE 3.0 Technical Report [research2024exaone3078binstruction] and the LG AI Ethics Principles [lgethics].
@@ 3.1 Benchmarks
  • World Knowledge We select benchmarks to evaluate the extent of our model’s world knowledge, including MMLU-Redux [gema2025mmlu] and MMLU-Pro [wang2024mmluprorobustchallengingmultitask], a refined and extended version of MMLU [hendrycks2021measuringmassivemultitasklanguage]. Additionally, we utilize GPQA-Diamond [rein2023gpqagraduatelevelgoogleproofqa] to assess the expert-level knowledge in Biology, Physics, and Chemistry.
@@
  • Math/Coding Challenging benchmarks in Math and Coding categories are adopted to evaluate the test-time computational capability of EXAONE 4.0. For Math, we utilize two math Olympiad competitions: AIME 2025 [maa2025aime] and HMMT Feb 2025 [balunovic_srimatharena_2025]. For Coding, LiveCodeBench v5 and v6 [jain2024livecodebenchholisticcontaminationfree] are chosen.
@@
  • Instruction Following To evaluate how well our models understand and align with users’ instructions, we select IFEval [zhou2023instructionfollowingevaluationlargelanguage] and Multi-IF [he2024multiifbenchmarkingllmsmultiturn], the latter being an extension of IFEval to support multi-turn and multilingual scenarios. We use only the English subset of Multi-IF to focus on assessing the multi-turn instruction-following ability on English.
@@
  • Long Context To evaluate the model’s ability to understand and solve tasks requiring long-context comprehension, we adopt three representative benchmarks: HELMET [yen2025helmetevaluatelongcontextlanguage], RULER [hsieh2024rulerwhatsrealcontext], and LongBench [bai2024longbenchbilingualmultitaskbenchmark]. These benchmarks collectively cover both synthetic tasks and real-world scenarios. To maintain a coherent evaluation focused on core long-context capabilities, we exclude the LongCite task from HELMET (see Appendix D for further details).
@@
  • Agentic Tool Use With the advancement of LLM-based agents, numerous benchmarks have emerged to evaluate their tool-use capabilities, among which we focus on the two most widely adopted: BFCL-v3 [patil2025bfcl] and Tau-Bench [yao2024taubenchbenchmarktoolagentuserinteraction]. BFCL-v3 evaluates various aspects of function-calling abilities. Tau-Bench assesses tool calling performance through simulated conversations between a user LLM. We utilize gpt-4.1-2025-04-14 model as the user role.
@@
  • Multilinguality Beyond English, we evaluate our models on two additional languages: Korean and Spanish. For Korean, we use KMMLU-Pro [^0] https://huggingface.co/datasets/LGAI-EXAONE/KMMLU-Pro for measuring practical applicability on professional knowledge and KMMLU-Redux [^1] https://huggingface.co/datasets/LGAI-EXAONE/KMMLU-Redux for assessing real-world expert knowledge instead of KMMLU [son-etal-2025-kmmlu] to ensure benchmark reliability. KMMLU have been reported dataset error and contamination issue between pre-training corpora and task dataset. In addition, we employ Korean School Math (KSM) subset of HRM8K [ko2025understandsolvetranslatebridging] to evaluate a wide range of Korean mathematical knowledge from high-school to Olympiad level. To evaluate the models’ ability to handle long-context Korean inputs, we also include an in-house benchmark, Ko-LongBench (Please refer to Appendix D.4 for details). For Spanish, we adopt the translated version of existing benchmarks. To be specific, we use MMMLU (es)  [^2] https://huggingface.co/datasets/openai/MMMLU and MATH500 [lightman2023lets] (es)  [^3] https://huggingface.co/datasets/bezir/MATH-500-multilingual. Furthermore, we assess translation ability using WMT24++ [deutsch2025wmt24expandinglanguagecoverage], a widely-used translation benchmark. We consider only EN and ES pair, and utilize LLM-as-a-judge 444gpt-4.1-2025-04-14 is used for the judge model. We follow reference-based direct assessment method used in WMT24++ [deutsch2025wmt24expandinglanguagecoverage]. The exact prompt used for judge is in Appendix D.5. to score the translation quality.
@@ 3.3 Experimental Setup
We sample $n$ different responses for each problem in benchmarks with limited examples to ensure evaluation stability. Specifically, we sample $n=8$ responses for GPQA-Diamond, $n=32$ for AIME 2025 and HMMT Feb 2025, and $n=4$ for LiveCodeBench v5/6, Tau-Bench and MATH500 (es). The accuracy is averaged over the $n$ samples. In Reasoning mode, we set temperature to 0.6, top-p [Holtzman2020The] to 0.95, and apply a presence penalty of 1.5 only for our 32B model. In contrast, for Non-Reasoning mode, greedy decoding is used for a single ($n=1$) generated response, while the same sampling settings as Reasoning mode (except with a presence penalty of 0.0) are used when generating $n>1$ responses. We generate a maximum of 64K tokens for AIME 2025, HMMT Feb 2025, LiveCodeBench v5/6, and KSM benchmarks, while 32K for other benchmarks.
@@
In evaluating long-context performance of Small-Size NON-REASONING models, we extend the context lengths of Qwen3 1.7B and Qwen3 0.6B beyond their 32K token limit by applying YaRN [peng2024yarn], enabling inference up to 64K tokens. For reference, evaluation results of models such as Gemma-3-1B, EXAONE-3.5-2.4B-Instruct, Qwen3 1.7B, and Qwen3 0.6B at context lengths up to 32K tokens are provided in the Appendix D.
@@ Appendix C Baseline models
Frontier DeepSeek R1-0528 [deepseekai2025deepseekr1incentivizingreasoningcapability] 671B (MoE) Reasoning May 2025
DeepSeek V3-0324 [deepseekai2025deepseekv3technicalreport] 671B (MoE) Non-reasoning Mar. 2025
@@
Qwen 3 235B [yang2025qwen3technicalreport] 235B (MoE) Hybrid Apr. 2025
Mid-size Qwen 3 32B [yang2025qwen3technicalreport] 32.8B Hybrid Apr. 2025
@@
Gemma 3 27B [gemmateam2025gemma3technicalreport] 27.4B Non-reasoning Mar. 2025
@@
Magistral-Small-2506 [mistralai2025magistral] 23.6B Reasoning Jun. 2025
Phi 4 reasoning plus [abdin2025phi4reasoningtechnicalreport] 14.7B Reasoning Apr. 2025
Phi 4 [abdin2024phi4technicalreport] 14.7B Non-reasoning Dec. 2024
@@
EXAONE Deep 2.4B [research2025exaonedeepreasoningenhanced] 2.41B Reasoning Mar. 2025
Qwen 3 1.7B [yang2025qwen3technicalreport] 1.72B Hybrid Apr. 2025
@@
Gemma 3 1B [gemmateam2025gemma3technicalreport] 1.00B Non-reasoning Mar. 2025
Qwen 3 0.6B [yang2025qwen3technicalreport] 596M Hybrid Apr. 2025